# Minimax surfaces and the holographic entropy cone

### v1

@@ 4.2.2 Spherically symmetric matter
then $\delta\phi=\phi_{\infty}(L,E)-\phi_{\infty}(L+\epsilon,-E)$ and $\delta t=t_{\infty}(L,E)-t_{\infty}(L+\epsilon,-E)$. Enforcing $\delta\phi\geq\delta t$ leads to a causality contraint $E\leq E_{c}$. For fixed $(L,\epsilon)$ not all values of $20<\delta<$ 2 are allowed to obtain a boosted H configuration; there exists a critical value of $\delta$ for which the two ribs meet at the same point on the boundary, and for higher values the two ribs cross before reaching the boundary thereby introducing new intersections between time-sheets, which is beyond the scope of this analysis. So when computing $E_{c}$, we filter out these cases. The overall trend we observe is for $E_{c}$ to be decreasing monotonically as a function of $\delta$, see fig 19.
@@
We show our results for (1) and (2) in figs. 20 and 21 respectively.
@@ 4.4 Geometric proof of the inequalities
As established in section 4.1, if the cooperating property holds, it implies the existence of a spacetime graph model which is graph-equivalent to the model for static states. It then follows directly that all holographic entropy inequalities valid for static states will remain valid in the time-dependent setting. However, we find instructive to spell out how these inequalities could be proven geometrically using minimax. In this section we assume the cooperating property and present a strategy which resembles the original exclusion/inclusion argument of Headrick:2007km ; Headrick:2013zda . We find this approach useful because it demonstrates how the minimax prescription utilizes the full spacetime (overcoming issues that maximin faced for higher inequalities) while also clarifying the role of the cooperating property.
@@
    The second relevant surface will be $\gamma^{\prime}(B)$ obtained as the union of the portions of $\gamma^{+}(BC)$ and $\gamma^{+}(AB)$ on $\tau^{\prime}(B)$ (i.e. the partial minimax surfaces which we refered to before in 3.4 and 4.1). Note that $\gamma^{\prime}(B)$ will generally be disconnected, as $\gamma^{+}(AB)$ and $\gamma^{+}(BC)$ need not intersect.
@@ 5 Discussion
While we hope to have given good evidence in favor of the cooperating property, it still remains a conjecture that needs further investigation. In particular, the comments in section 4.3 suggest that the conjecture may be false in its current form. However, we are hopeful that under some mild modifications a weaker statement may be proven that still allows for the construction of the graph model. If even that turns out to be false, we note that the cooperating property is only a sufficient condition for our results, and there is still overwhelming evidence in favor of the equivalence between the RT and HRT cones, as we argued in the introduction. This would make the search for such an elusive proof even more exciting.
@@ A.1 AdS3
where $V_{\text{eff}}=r^{2}+1+E^{2}-L^{2}-\frac{L^{2}}{r^{2}}$ is an effective potential. Substituting $u=r^{2}$ we have
$\frac{\dot{u}^{2}}{4}=-u^{2}-au+\operatorname{\mathsf{L}}^{2}$ (69)
@@
$t(\lambda)=\tan^{-1}\left[\left(\frac{-2-a+\sqrt{\Delta}}{2E}\right)\tanh\lambda\right],$ (73)
@@
$\phi(\lambda)=\tan^{-1}\left[\left(\frac{-a+\sqrt{\Delta}}{2L}\right)\tanh\lambda\right]$ (74)
It’s useful to rewrite the above formulas not in terms of $(E,L)$ but in terms of the opening angle $\Delta\phi$ and the time separation $\Delta t$ between a region’s endpoints. These are easily found by integrating eqs 66. One finds
$\Delta t=2\tan^{-1}\left(\frac{-2-a+\sqrt{\Delta}}{2E}\right),\quad\Delta\phi=2\tan^{-1}\left(\frac{-a+\sqrt{\Delta}}{2L}\right).$ (75)
@@
$s=(0,1,0),\quad r=\left(\frac{-2-a+\sqrt{\Delta}}{2E},0,\frac{-a+\sqrt{\Delta}}{2L}\right)=\left(\tan{\frac{\Delta t}{2}},0,\tan{\frac{\Delta\phi}{2}}\right)$ ) ) (80)

### v2

@@ 4.2.2 Spherically symmetric matter
then $\delta\phi=\phi_{\infty}(L,E)-\phi_{\infty}(L+\epsilon,-E)$ and $\delta t=t_{\infty}(L,E)-t_{\infty}(L+\epsilon,-E)$. Enforcing $\delta\phi\geq\delta t$ leads to a causality contraint $E\leq E_{c}$. For fixed $(L,\epsilon)$ not all values of $20<\delta<$ 2 are allowed to obtain a boosted H configuration; there exists a critical value of $\delta$ for which the two ribs meet at the same point on the boundary, and for higher values the two ribs cross before reaching the boundary thereby introducing new intersections between time-sheets, which is beyond the scope of this analysis. So when computing $E_{c}$, we filter out these cases. The overall trend we observe is for $E_{c}$ to be decreasing monotonically as a function of $\delta$, see figure 19.
@@
We show our results for (1) and (2) in figures 20 and 21 respectively.
@@ 4.4 Geometric proof of the inequalities
As established in subsection 4.1, if the cooperating property holds, it implies the existence of a spacetime graph model which is graph-equivalent to the model for static states. It then follows directly that all holographic entropy inequalities valid for static states will remain valid in the time-dependent setting. However, we find instructive to spell out how these inequalities could be proven geometrically using minimax. In this section we assume the cooperating property and present a strategy which resembles the original exclusion/inclusion argument of Headrick:2007km ; Headrick:2013zda . We find this approach useful because it demonstrates how the minimax prescription utilizes the full spacetime (overcoming issues that maximin faced for higher inequalities) while also clarifying the role of the cooperating property.
@@
    The second relevant surface will be $\gamma^{\prime}(B)$ obtained as the union of the portions of $\gamma^{+}(BC)$ and $\gamma^{+}(AB)$ on $\tau^{\prime}(B)$ (i.e. the partial minimax surfaces which we refered to before in subsections 3.4 and 4.1). Note that $\gamma^{\prime}(B)$ will generally be disconnected, as $\gamma^{+}(AB)$ and $\gamma^{+}(BC)$ need not intersect.
@@ 5 Discussion
While we hope to have given good evidence in favor of the cooperating property, it still remains a conjecture that needs further investigation. In particular, the comments in subsection 4.3 suggest that the conjecture may be false in its current form. However, we are hopeful that under some mild modifications a weaker statement may be proven that still allows for the construction of the graph model. If even that turns out to be false, we note that the cooperating property is only a sufficient condition for our results, and there is still overwhelming evidence in favor of the equivalence between the RT and HRT cones, as we argued in the introduction. This would make the search for such an elusive proof even more exciting.
@@ A.1 AdS3
where $V_{\text{eff}}=-r^{2}-1-E^{2}+L^{2}+\frac{L^{2}}{r^{2}}$ is an effective potential. Substituting $u=r^{2}$ we have
$\frac{\dot{u}^{2}}{4}=u^{2}+au-L^{2}$ (71)
@@
$t(\lambda)=\tan^{-1}\left[\left(\frac{-2+a+\sqrt{\Delta}}{2E}\right)\tanh\lambda\right],$ (75)
@@
$\phi(\lambda)=\tan^{-1}\left[\left(\frac{a+\sqrt{\Delta}}{2L}\right)\tanh\lambda\right]$ (76)
It’s useful to rewrite the above formulas not in terms of $(E,L)$ but in terms of the opening angle $\Delta\phi$ and the time separation $\Delta t$ between a region’s endpoints. These are easily found by integrating equations (68). One finds
$\Delta t=2\tan^{-1}\left(\frac{-2+a+\sqrt{\Delta}}{2E}\right),\quad\Delta\phi=2\tan^{-1}\left(\frac{a+\sqrt{\Delta}}{2L}\right).$ (77)
@@
$s=(0,1,0),\quad r=\left(\frac{-2+a+\sqrt{\Delta}}{2E},0,\frac{a+\sqrt{\Delta}}{2L}\right)=\left(\tan{\frac{\Delta t}{2}},0,\tan{\frac{\Delta\phi}{2}}\right)$ ) ) (82)